A +100 μC charge is placed at the origin. A -50 μC is placed at x=2m and a +200 μC is placed at x= -4m.

What is the net electric force acting on the +100 μC charge?





1
Expert's answer
2021-04-05T11:08:09-0400

Let "q_1 = +100\mu C = 0.1\times 10^{-3}C, q_2 = -0.05\times 10^{-3}C, q_3 = 0.2\times 10^{-3}C"

The distance between "q_1" and "q_2" is "r_{12} = 2m", the distance between "q_1" and "q_3" is "r_{13} = 4m".

Then, according to the Coulomb's law, the force on "q_1" from "q_2" is:



"F_{12} =k \dfrac{q_1q_2}{r_{12}^2}"

where "k\approx 9\times 10^9N\cdot m^2/C^2" is the Coulomb's constant.


"F_{12} =9\times 10^9\cdot \dfrac{0.1\times 10^{-3}\cdot (-0.05\times 10^{-3})}{2^2} = -11.25N"

The force on "q_1" from "q_3" is:


"F_{13} =k \dfrac{q_1q_3}{r_{13}^2}""F_{13} =9\times 10^9\cdot \dfrac{0.1\times 10^{-3}\cdot 0.2\times 10^{-3}}{2^2} = 45N"


Since these forces are directed along one line, we can add them. The resultant force is then:


"F_1 = F_{12} + F_{13} = -11.25N + 45N = 33.75N"

And directed as shown in figure below:




Answer. 33.75 N.